OpenStudy (anonymous):

The equation of a line with slope m = -3 and including point (5, 5) is y = -3x + 20. true or false

OpenStudy (danjs):

the equation is in form y = m*x+b does the M value match

OpenStudy (danjs):

and put in x = 5 and y=5 , do both sides equal?

OpenStudy (calculusxy):

@DanJS really said what you should do. since u know that m = -3 (slope) and the slope-intercept form is: \[y = mx + b \] and your coordinates are: \[(x,y) = (5,5)\] so now you just need to replace 5 for x and 5 for y. \[5 = -3(5) + 20\] when you compute\[-3(5) + 20\] do you get 5? if yes, then it's true. if not, then it's false. *sorry if you did exactly what i did in your attachment, but i couldn't open it up, so i wrote this whole thing ;) *

OpenStudy (calculusxy):

@kacrabtree sorry, nvm about the attachment thing :(